OpenStudy (anonymous):

SOME ONE PLEASE HELP! •If you chose translation, use the coordinates of your transformation along with the distance formula to show that the two triangles are congruent by the SSS postulate. You must show all work with the distance formula and each corresponding pair of sides to receive full credit. I am so confused

OpenStudy (danjs):

i can't view the picture, but use the distance formula for 2 points to show that the translated side is the same as before, for all 3 sides

OpenStudy (danjs):

\[distance = \sqrt{(x2 - x1)^2 + (y2 - y1)^2}\]

OpenStudy (anonymous):

@DanJS sorry can you see it now?

OpenStudy (danjs):

ok, translating that does not change its size, just moves it around the coordinate plane, the sides of the translated triangle should be the same lengths as the sides of the original

OpenStudy (anonymous):

so I use the equation you said to solve?

OpenStudy (danjs):

the vertical and horizontal lines, you can just measure lengths directly, vertical 3 to -3 = 6 units horizontal -2 to 6 = 8 units use that distance formula to calculate the diagonal length

OpenStudy (danjs):

(x1,y1) =(-2, -3) to (x2. y2)=(6,3)

OpenStudy (anonymous):

im so sorry I am so bad at math I am trying to understand

OpenStudy (danjs):

\[distance = \sqrt{(x2 - x1)^2 + (y2 - y1)^2}\] \[distance = \sqrt{(6 - (-2))^2 + (3 - (-3))^2} = \sqrt{8^2+6^2}\]

OpenStudy (anonymous):

I GET IT! THANK YOU!

OpenStudy (danjs):

you are just using the pythagorean theorem to calculate the hypotenuse length for a right triangle c^2 = a^2 + b^2 where c is the diagonal, a and b are the side lengths

OpenStudy (anonymous):

So then I just solve that equation and that's the distance between those points?

OpenStudy (danjs):

yep, then translate it somewhere else, and the sides should be the same length as before, showing SSS congruent triangles

OpenStudy (anonymous):

so to translate I just draw the triangle again?

OpenStudy (danjs):

i guess, choose translate and move the triangle i think, each point changes the same amount in any direction so that the moved triangle should still have all lengths the same for each side,

OpenStudy (anonymous):

oh okay! thank you so much I really appreciate it! how do I give you a medal?

OpenStudy (danjs):

like move all 3 points over 3 and up 2 for example,

OpenStudy (danjs):

welcome, meh medals

OpenStudy (anonymous):

thank you !!!!

OpenStudy (anonymous):

@DanJS I am so sorry to bother you again but, I got the distance equals 10. Now I am confused on what they want me to do next.

OpenStudy (danjs):

translate the triangle , not sure what the program has you do, move all 3 points by the same (x + a . y +b), you pick 'a' and 'b'

OpenStudy (danjs):

the resulting triangle will have a new location on the grid, but will be the exact same as the old triangle in terms of side lengths, use the calculation you did before to get 10, but this time with the translated points

OpenStudy (anonymous):

oh okay so when I use the translated points I should get 10 again? then I am done?

OpenStudy (danjs):

yeah, you have to show all 3 sides have the same lenghts as before

OpenStudy (danjs):

you are just sliding the thing , so of course the side lengths will be the same right?

OpenStudy (danjs):

not squishing it or dilation , where the sides will change

OpenStudy (anonymous):

exactly! thank you again!:)